# Generalized Wigner–Smith theory for perturbations at exceptional and diabolic point degeneracies

### v1

@@ Abstract
Spectral degeneracies, including diabolic (DP) and exceptional (EP) points, exhibit unique sensitivity to external perturbations, enabling powerful control and engineering of wave phenomena. We present a residue-based perturbation theory that quantifies complex resonance splitting of DP and EP type spectral degeneracies using generalized Wigner–Smith operators. We validate our theory using both analytic Hamiltonian models and numerical electromagnetic simulations, demonstrating excellent agreement across a range of cases. Our approach accurately predicts degenerate resonance splitting using only scattering data, offering a powerful framework for precision tuning, inverse design, and practical exploitation of non-Hermitian phenomena.
@@
Introduction–The Wigner–Smith (WS) time delay operator, initially introduced in quantum scattering theory Wigner (1955); Smith (1960), serves as a fundamental tool for characterizing the temporal response of resonant systems. For a system described by a scattering matrix $\mathbf{S}(\omega)$, the eigenvalues of the WS matrix $\mathbf{Q}_{\omega}=-i\,\mathbf{S}^{-1}\ \partial_{\omega}\mathbf{S}$ quantify the group delays of a set of incident wavepackets with modal structures given by the corresponding eigenvectors Patel and Michielssen (2021). $\mathbf{Q}_{\omega}$ connects a system’s frequency-domain response to its energetic and temporal properties, allowing studies of, e.g., time-delay statistics in chaotic cavities Grabsch et al. (2018), light storage in random media Durand et al. (2019), and density of states enhancement in nuclear physics Higgins et al. (2020). More recently, generalized WS (GWS) operators of the form $\mathbf{Q}_{\xi}=-i\,\mathbf{S}^{-1}\partial_{\xi}\mathbf{S}$, where $\xi$ denotes an arbitrary system parameter, have been proposed, extending conventional WS operators into parameter space and connecting external scattering observables to internal parametric sensitivities Horodynski et al. (2020). A key feature of GWS operators is that their principal modes, defined in analogy to the dispersion-free principal modes derived from $\mathbf{Q}_{\omega}$ Carpenter et al. (2015), have been shown to exhibit insensitivity to the variable conjugate to $\xi$ Ambichl et al. (2017). This insight has lead to a wealth of applications and recent advances in wavefront shaping and control Del Hougne et al. (2021), including targeted focusing in complex media Sol et al. (2025), guided light delivery in deformed optical fibers Matthès et al. (2021), manipulation of optical forces and trap stiffness in microparticle systems Būtaitė et al. (2024), tailored energy redistribution in time-varying systems Hüpfl et al. (2023), and the construction of optimal information states for scattering measurements Bouchet et al. (2021).
@@
Recent studies have extended GWS operators to describe parametric spectral shifts in non-Hermitian systems Byrnes and Foreman (2025a, b). In particular, Byrnes and Foreman Byrnes and Foreman (2025a) demonstrated that a complex analytic treatment of the GWS operator yields compact perturbative formulae for shifts in isolated resonances or anti-resonances associated with the poles and zeros of $\mathbf{S}$. Their treatment, however, was restricted to non-degenerate resonances and therefore fails at diabolic points (DPs), where multiple modes share the same frequency Berry and Wilkinson (1984), and exceptional points (EPs), where degenerate modes also become linearly dependent Miri et al. (2019). These degeneracies are not only mathematically rich, but also practically relevant. For example, DPs underpin mode-splitting sensors Zhu et al. (2010), while EPs enable enhanced sensitivity, chiral mode control, and topological transport Lai et al. (2019); Zhang et al. (2019); Ding et al. (2022). A GWS framework that rigorously handles such spectral singularities remains lacking. In this work, we address this gap by developing a residue-based formalism that extends earlier perturbative formulae to arbitrary-order DP and EP type spectral degeneracies and naturally reduces to earlier results in the non-degenerate case. We validate our derived formula using both analytic Hamiltonian models and electromagnetic simulations, thereby demonstrating its applicability to nanophotonic structures.
@@
GWS perturbation theory at degenerate poles–We begin by specifying the types of systems under study. Consider a physical system with a non-Hermitian $N\times N$ Hamiltonian matrix $\mathbf{H}(\alpha)\in\mathbb{C}^{N\times N}$ depending on a complex control parameter $\alpha$. We assume that for a particular parameter value $\alpha=\alpha_{0}$, $\mathbf{H}(\alpha_{0})$ has a repeated eigenvalue $\omega_{p}$ with algebraic multiplicity $\mathrm{AM}=N$. The nature of the degeneracy and the behavior of the system in response to perturbations is determined by the geometric multiplicity, $\mathrm{GM}$, which equals the number of distinct Jordan blocks associated with $\omega_{p}$. When $\mathrm{AM}=\mathrm{GM}$, $\mathbf{H}(\alpha_{0})$ is diagonalizable and the degeneracy corresponds to a DP. When $\mathrm{GM}<\mathrm{AM}$, however, $\mathbf{H}(\alpha_{0})$ is defective and contains at least one Jordan block of size greater than one. For our purposes, an EP will correspond to the case $\mathrm{GM}=1$, meaning $\mathbf{H}(\alpha_{0})$ contains a single Jordan block of size $N\times N$. Intermediate cases $1<\mathrm{GM}<N$ correspond to hybrid DP-EP scenarios and shall be discussed briefly below.
@@
where $\Delta\alpha=\alpha-\alpha_{0}$ and $c\in\mathbb{C}\neq 0$. We note immediately that Eq. (1) implies that, to leading order, the degenerate eigenvalue splits symmetrically and shifts in proportion to $\Delta\alpha^{1/N}$. Deviations from this scaling have been reported for systems experiencing non-generic perturbations, such as those that preserve symmetry Grom et al. (2025). Such effects, however, are beyond the scope of this work.
In addition to $\Delta\alpha$, the shift sensitivity is also governed by the coefficient $c$, which remains to be determined. It has been shown elsewhere that this coefficient is related to the orthogonality of the eigenvectors of $\mathbf{H}(\alpha_{0})$ Wiersig (2023). Here we show that it can also be related to the scattering matrix $\mathbf{S}\in\mathbb{C}^{M\times M}$, defined by Mahaux and Weidenmüller (1969)
@@
where $\mathbf{I}_{M}$ ($\mathbf{I}_{N}$) is the $M\times M$ ($N\times N$) identity matrix and $\mathbf{W}\in\mathbb{C}^{N\times M}$ couples the $N$ internal modes of the system to $M$ external scattering channels. To achieve this, we use the result Welters (2011)
$c^{N}=-\left.\frac{\partial}{\partial\alpha}\,\det\bigl{[}\omega\mathbf{I}_{N}-\mathbf{H}(\alpha)\bigr{]}\right|_{\begin{subarray}{c}\omega=\omega_{p}\\ \alpha=\alpha_{0}\end{subarray}}$ (3)
in conjunction with Sylvester’s identity Akritas et al. (1996), the latter of which can be used to show that
@@
where $g(\omega,\alpha)=\det\bigl{[}\omega\mathbf{I}_{N}-\mathbf{H}(\alpha)-i\mathbf{W}\mathbf{W}^{\dagger}\bigr{]}$. Differentiating (4) with respect to $\alpha$ and applying Jacobi’s identity to the derivative of $\det(\mathbf{S}^{-1})$ yields
@@
where we have used $\det[\omega\mathbf{I}_{N}-\mathbf{H}(\alpha_{0})]=(\omega-\omega_{p})^{N}$. To proceed, we assume that $\mathbf{W}$ is such that $g$ is analytic and non-zero at the EP, which implies the second term in (5) vanishes. Though this condition is not always satisfied, it is readily met when $\mathbf{W}$ is full rank. In practice, this condition is desirable because it eliminates the presence of dark modes. Standard limit laws allow us to write $c=\mathrm{Res}_{\omega=\omega_{p}}\{[i\,\mathrm{tr}\,\mathbf{Q}_{\alpha}(\omega,\alpha_{0})]^{1/N}\}$, which can be substituted back into (1). Neglecting higher order terms then yields the resonance shift formula
@@
One simple class of $\mathbf{W}$ matrices for which the eigenvalues of $\mathbf{H}$ do not mix among those of $\mathbf{S}$ can be found by observing the algebraic structure of Eq. (2). Note first that although $\mathbf{W}$ is not in general square, it possess a right inverse $\mathbf{W}_{R}=\mathbf{W}^{\dagger}(\mathbf{W}\mathbf{W}^{\dagger})^{-1}$. If $\mathbf{W}\mathbf{W}^{\dagger}=\gamma\mathbf{I}_{N}$ for some constant $\gamma$, then it follows that $\mathbf{W}^{\dagger}=\gamma\mathbf{W}_{R}$ and Eq. (2) thus expresses pseudo-similarity between $\mathbf{S}$ and $\mathbf{H}$. Physically, this condition dictates that each internal mode couples to the exterior through orthogonal, non-interfering channels with uniform coupling strength $\gamma$. The eigenvalues of $\mathbf{S}$ will then be of the form $\mu_{n}=1-i\gamma(\omega-\omega_{n})^{-1}$ and analysis of the operators $Q_{\alpha,n}=-i\mu_{n}^{-1}\partial_{\alpha}\mu_{n}$ yields the pole shifts. Though this condition is sufficient to avoid mixing, it is not necessary, and pole mixing can be avoided even in cases where $\mathbf{W}\mathbf{W}^{\dagger}$ is not diagonal. An exhaustive analysis of such cases is beyond the scope of this paper, but we note that if the form of $\mathbf{W}$ is known, one can always compute its inverse directly and obtain the resolvent $(\omega\mathbf{I}_{N}-\mathbf{H})^{-1}=-i\mathbf{W}_{R}^{\dagger}(\mathbf{S}-\mathbf{I}_{M})\mathbf{W}_{R}$, whose eigenvalues each are of the form $(\omega-\omega_{n})^{-1}$. This calculation reverts the problematic mixing caused by $\mathbf{W}$, yielding a set of functions that can be used to find the pole shifts. As a final remark, we note that the cases $1<\mathrm{GM}<N$ present similar challenges associated with the structure of $\mathbf{W}$. In principle, however, the analysis can be restricted to each Jordan block individually, allowing the techniques described here to be applied separately to each one.
@@
DP in polarizability tensor of ellipsoidal nanoparticles–We now consider the case of a DP in a small, homogeneous, metallic ellipsoid. Our ellipsoid, embedded in a background of unity permittivity, has semi-axes $a_{1},a_{2},a_{3}$ (see inset of Figure 1) and complex permittivity $\varepsilon(\omega)$ taken as that of gold assuming a Drude–Lorentz model Rakić et al. (1998). In a coordinate system aligned with the ellipsoid’s principal axes, the $3\times 3$ polarizability tensor is diagonal with entries Bohren and Huffman (2008)
@@
Figure 2: Splitting in the complex plane of (a) EP2, (b) EP3, and (c) EP4 degenerate resonances supported in TO designed plasmonic nanowire structures (insets) upon a perturbation, $\Delta\varepsilon$, of the host permittivity. Complex frequency shifts were calculated from direct solution of the TO dispersion relation ($\circ$ markers) and our GWS residue formula ($\times$ markers). Marker colour encodes $\Delta\varepsilon$ on a logarithmic scale. (d) Log–log plot of the magnitude of the frequency shift $|\Delta\omega|$ versus $\Delta\varepsilon$ for each EP. Gradients of $1/2$, $1/3$, and $1/4$ for EP2, EP3 and EP4 respectively (black dashed fits) confirm the expected $\Delta\varepsilon^{1/N}$ scaling.
@@
In conclusion, we have developed a perturbative framework that extends prior GWS perturbation theory to non-Hermitian systems exhibiting EPs and DPs. Our approach shares the merits of its previous iterations, enabling the determination of internal resonant behavior from external scattering measurements in a flexible and straightforward way. Our theory was validated through analytic Hamiltonian models and electromagnetic simulations of assorted nanophotonic and TO engineered plasmonic systems. Our results demonstrate the method’s accuracy, generality, and computational practicality, offering new perspectives and tools for resonance control, precision sensing, and inverse design of complex non-Hermitian environments.
@@ Acknowledgements.
K.W. was supported by a Nanyang Technological University Interdisciplinary Graduate Program (NTU-IGP) Research Scholarship. N.B. was supported by Singapore Ministry of Education Academic Research Fund (Tier 1) Grant RG66/23. M.R.F. was supported by funding from the Institute for Digital Molecular Analytics and Science (IDMxS) under the Singapore Ministry of Education Research Centres of Excellence scheme (EDUN C-33-18-279-V12) and by Nanyang Technological University Grant SUG:022824-00001.

### v2

@@ Abstract
Resonant spectral degeneracies, including diabolic (DP) and exceptional (EP) points, exhibit unique sensitivity to external perturbations, enabling powerful control and engineering of wave phenomena. We present a residue-based perturbation theory that quantifies complex resonance splitting of DP and EP type spectral degeneracies using generalized Wigner–Smith operators. We validate our theory using both analytic Hamiltonian models and numerical electromagnetic simulations, demonstrating excellent agreement across a range of cases. Our approach accurately predicts degenerate resonance splitting using only scattering data, offering a powerful framework for sensitivity analysis, inverse design, and quantitative sensing.
@@
The Wigner–Smith (WS) time delay operator, initially introduced in quantum scattering theory Wigner (1955); Smith (1960), serves as a fundamental tool for characterizing the temporal response of resonant systems. For a system described by a scattering matrix $\mathbf{S}(\omega)$, the real part of the eigenvalues of the WS matrix $\mathbf{Q}_{\omega}=-i\,\mathbf{S}^{-1}\ \partial_{\omega}\mathbf{S}$ quantify the group delays of a set of incident wavepackets with modal structures given by the corresponding eigenvectors Patel and Michielssen (2021). $\mathbf{Q}_{\omega}$ connects a system’s frequency-domain response to its energetic and temporal properties, allowing studies of, e.g., time-delay statistics in chaotic cavities Grabsch et al. (2018), light storage in random media Durand et al. (2019), and density of states enhancement in nuclear physics Higgins et al. (2020). More recently, generalized WS (GWS) operators of the form $\mathbf{Q}_{\xi}=-i\,\mathbf{S}^{-1}\partial_{\xi}\mathbf{S}$, where $\xi$ denotes an arbitrary system parameter, have been proposed, extending conventional WS operators into parameter space and connecting external scattering observables to internal parametric sensitivities Horodynski et al. (2020). A key feature of GWS operators is that their principal modes, defined in analogy to the dispersion-free principal modes derived from $\mathbf{Q}_{\omega}$ Carpenter et al. (2015), have been shown to exhibit insensitivity to the variable conjugate to $\xi$ Ambichl et al. (2017). This insight has led to a wealth of advances in wavefront shaping and control Del Hougne et al. (2021), including targeted focusing in complex media Sol et al. (2025), guided light delivery in deformed optical fibers Matthès et al. (2021), manipulation of optical forces and trap stiffness in microparticle systems Būtaitė et al. (2024), tailored energy redistribution in time-varying systems Hüpfl et al. (2023), and the construction of optimal information states for scattering measurements Bouchet et al. (2021).
@@
Recent studies have extended GWS operators to describe parametric spectral shifts in non-Hermitian systems Byrnes and Foreman (2025a, b). In particular, Byrnes and Foreman Byrnes and Foreman (2025a) demonstrated that a complex analytic treatment of the GWS operator yields compact perturbative formulae for shifts in isolated resonances or anti-resonances associated with the poles and zeros of $\mathbf{S}$. Their treatment, however, was restricted to non-degenerate resonances associated with simple poles of $\mathbf{S}$ and therefore fails at diabolic points (DPs), where multiple resonant modes share the same frequency Berry and Wilkinson (1984), and exceptional points (EPs), where resonant modes degenerate in frequency also become linearly dependent Miri et al. (2019). These degeneracies are not only mathematically rich, but also practically relevant. For example, DPs underpin mode-splitting sensors Zhu et al. (2010), while EPs enable chiral mode control, topological transport, and enhanced sensitivity Lai et al. (2019); Li et al. (2023); Loughlin and Sudhir (2024). A GWS framework that rigorously handles degenerate resonances remains lacking. In this work, we address this gap by developing a residue-based formalism that extends earlier perturbative formulae to arbitrary-order DP and EP type degeneracies and naturally reduces to earlier results in the non-degenerate case. We validate our derived formula using both analytic Hamiltonian models and electromagnetic simulations, thereby demonstrating its applicability to nanophotonic structures.
@@
GWS perturbation theory at degenerate poles– We begin by briefly reviewing the key concepts from non-Hermitian physics most relevant to our study. More information can be found in the supplemental material and the literature Ashida et al. (2020). Consider a physical system with a non-Hermitian $N\times N$ Hamiltonian matrix $\mathbf{H}(\alpha)\in\mathbb{C}^{N\times N}$ depending on a complex control parameter $\alpha$. Our work concerns resonant DP and EPs associated with the eigenmodes of $\mathbf{H}$, as opposed to scattering DP and EPs associated with those of $\mathbf{S}$ Sweeney et al. (2019). We assume that for a particular parameter value $\alpha=\alpha_{0}$, $\mathbf{H}(\alpha_{0})$ has a repeated (degenerate) eigenvalue $\omega_{p}$ with algebraic multiplicity $\mathrm{AM}=N$ and geometric multiplicity $\mathrm{GM}$ satisfying $1\leq\mathrm{GM}\leq N$. When $\mathrm{AM}=\mathrm{GM}$, $\mathbf{H}(\alpha_{0})$ is diagonalizable and the degeneracy corresponds to a DP. When $\mathrm{GM}<\mathrm{AM}$, however, $\mathbf{H}(\alpha_{0})$ is defective and contains at least one Jordan block of size greater than one. For our purposes, an EP will correspond to the case $\mathrm{GM}=1$, meaning $\mathbf{H}(\alpha_{0})$ contains a single Jordan block of size $N\times N$. Intermediate cases $1<\mathrm{GM}<N$ correspond to hybrid DP-EP scenarios as discussed below.
@@
where $\Delta\alpha=\alpha-\alpha_{0}$ and $c\in\mathbb{C}\neq 0$. We note immediately that Eq. (1) implies that, to leading order, the degenerate eigenvalue splits symmetrically and shifts in proportion to $\Delta\alpha^{1/N}$. Deviations from this scaling have been reported for systems experiencing non-generic perturbations, such as those preserving symmetry Grom et al. (2025). Such cases are beyond the scope of this work.
In addition to $\Delta\alpha$, the shift sensitivity is also governed by the coefficient $c$, which remains to be determined. It has been shown elsewhere that this coefficient is related to the orthogonality of the eigenvectors of $\mathbf{H}(\alpha_{0})$ Wiersig (2023). Here we show that it can also be related to the scattering matrix $\mathbf{S}\in\mathbb{C}^{M\times M}$. A commonly used expression for $\mathbf{S}$ in terms of $\mathbf{H}$ is given by Fyodorov and Sommers (1997)
@@
where $\mathbf{I}_{M}$ is the $M\times M$ ($N\times N$) identity matrix and $\mathbf{W}\in\mathbb{C}^{N\times M}$ couples the $N$ internal modes of the system to $M$ external scattering channels. A key property of Eq. (2) is that degenerate eigenfrequencies of $\mathbf{H}$ are mapped to degenerate poles of $\mathbf{S}$ via the resolvent $(\omega\mathbf{I}_{N}-\mathbf{H}(\alpha))^{-1}$. It has also been shown that Welters (2011)
$c^{N}=-\left.\frac{\partial}{\partial\alpha}\,\det\bigl[\omega\mathbf{I}_{N}-\mathbf{H}(\alpha)\bigr]\right|_{{\omega=\omega_{p},~\alpha=\alpha_{0}}},$ (3)
which can be linked to Eq. (2) using the identity
@@
where $g(\omega,\alpha)=\det\bigl[\omega\mathbf{I}_{N}-\mathbf{H}(\alpha)-i\mathbf{W}\mathbf{W}^{\dagger}\bigr]$. A short derivation of Eq. (4) is given in the supplemental material. Differentiating (4) with respect to $\alpha$ and applying Jacobi’s identity to the derivative of $\det(\mathbf{S}^{-1})$, we can obtain $c^{N}$ by evaluating the limit
@@
where we have used $\det[\omega\mathbf{I}_{N}-\mathbf{H}(\alpha_{0})]=(\omega-\omega_{p})^{N}$. To proceed, we note that the final term in Eq. (5) vanishes provided that $\mathbf{W}$ is non-zero (see supplemental material). Standard limit laws allow us to write $c=\mathrm{Res}_{\omega=\omega_{p}}\{[i\,\mathrm{tr}\,\mathbf{Q}_{\alpha}(\omega,\alpha_{0})]^{1/N}\}$, which can be substituted back into (1). Neglecting higher order terms yields the resonance shift formula
@@
One simple class of $\mathbf{W}$ matrices for which the eigenvalues of $\mathbf{H}$ do not mix among those of $\mathbf{S}$ can be found by observing the algebraic structure of Eq. (2). Note first that although $\mathbf{W}$ is not in general square, it possess a right inverse $\mathbf{W}_{R}=\mathbf{W}^{\dagger}(\mathbf{W}\mathbf{W}^{\dagger})^{-1}$. If $\mathbf{W}\mathbf{W}^{\dagger}=\gamma\mathbf{I}_{N}$ for some constant $\gamma$, then it follows that $\mathbf{W}^{\dagger}=\gamma\mathbf{W}_{R}$ and Eq. (2) thus expresses pseudo-similarity between $\mathbf{S}$ and $\mathbf{H}$. Physically, this condition dictates that each internal mode couples to the exterior through orthogonal, non-interfering channels with uniform coupling strength $\gamma$. The eigenvalues of $\mathbf{S}$ will then be of the form $\mu_{n}=1-i\gamma(\omega-\omega_{n})^{-1}$ and analysis of the operators $Q_{\alpha,n}=-i\mu_{n}^{-1}\partial_{\alpha}\mu_{n}$ yields the pole shifts. Though this condition is sufficient to avoid mixing, it is not necessary, and pole mixing can be avoided even in cases where $\mathbf{W}\mathbf{W}^{\dagger}$ is not diagonal. An example of such a case is presented in the supplemental material, but an exhaustive analysis is beyond the scope of this paper. In any case, we note that if the form of $\mathbf{W}$ is known, one can always compute its inverse directly and obtain the resolvent $(\omega\mathbf{I}_{N}-\mathbf{H})^{-1}=-i\mathbf{W}_{R}^{\dagger}(\mathbf{S}-\mathbf{I}_{M})\mathbf{W}_{R}$, whose eigenvalues each are of the form $(\omega-\omega_{n})^{-1}$. This calculation reverts the problematic mixing caused by $\mathbf{W}$, yielding a set of functions that can be used to find the pole shifts. As a final remark, we note that the cases $1<\mathrm{GM}<N$ present similar challenges associated with the structure of $\mathbf{W}$. In principle, however, the analysis can be restricted to each Jordan block individually, allowing the techniques described to be applied to each one.
@@
DP in ellipsoidal nanoparticles–We now consider the case of a DP in a small, homogeneous, metallic ellipsoid. Our ellipsoid, embedded in a background of unity permittivity, has semi-axes $a_{1},a_{2},a_{3}$ (see inset of Figure 1) and complex permittivity $\varepsilon(\omega)$ taken as that of gold assuming a Drude–Lorentz model Rakić et al. (1998). In a coordinate system aligned with the ellipsoid’s principal axes, the $3\times 3$ polarizability tensor is diagonal with entries Bohren and Huffman (2008)
@@
Figure 2: Splitting in the complex plane of (a) EP2, (b) EP3, and (c) EP4 degenerate resonances supported in TO designed plasmonic nanowire structures (insets) upon a perturbation, $\Delta\varepsilon$, of the host permittivity. Complex frequency shifts were calculated from direct solution of the TO dispersion relation ($\circ$ markers) and our GWS residue formula ($\times$ markers). Marker colour encodes $\Delta\varepsilon$ on a logarithmic scale. (d) Log–log plot of the magnitude of the frequency shift $|\Delta\omega|$ versus $\Delta\varepsilon$ for each EP. Gradients of $1/2$, $1/3$, and $1/4$ are found for EP2, EP3 and EP4 respectively (black dashed fits).
In conclusion, we have developed a perturbative framework that extends prior GWS perturbation theory to non-Hermitian systems exhibiting EPs and DPs. Our approach shares the merits of previous iterations and enables determination of internal resonant behavior from external scattering measurements in a flexible and straightforward way Byrnes and Foreman (2025a, b). Our theory was validated through analytic modelling and electromagnetic simulations of assorted nanophotonic and TO engineered plasmonic systems. The final example in particular also demonstrates that although Eq. (6) was derived based on Eq. (2), it does not depend on this exact relation and can also be used in situations where neither $\mathbf{H}$ nor $\mathbf{S}$ are convenient to describe. Our results demonstrate the method’s accuracy, generality, and computational practicality, offering new perspectives and tools for sensitivity analysis, quantitative sensing, resonance control and inverse design of complex non-Hermitian environments (see e.g. Byrnes and Foreman (2025a, b)).
@@ Acknowledgements.
K.W. and N.B were funded by a Nanyang Technological University Interdisciplinary Graduate Program (NTU-IGP) Scholarship and Singapore Ministry of Education Academic Research Fund (Tier 1) Grant RG66/23 respectively. M.R.F. was supported by IDMxS under the Singapore Ministry of Education Research Centres of Excellence scheme (EDUN C-33-18-279-V12) and by NTU Grant SUG:022824-00001.